I tried to figure out where additional (to seismic) noise enters to MC_F, so the coherence below 1 Hz is low (~0.2-0.5). I've examined noise in the path
PD -> DEMOD -> MC BOARD -> FSS -> PZT Controller -> LASER
MC BOARD noise shows up only below 0.1 Hz and at 10 Hz where the whitening filter starts to work. SNR is ~2 at 4 Hz, so we might want to slightly improve whitening filter. But other then that path PD->DEMOD->MC BOARD is not responsible for additional noises below 1 Hz.
Next I connected SR785 to the laser and measured the closed loop feedback signal while MC was locked.
Coherence between MC BOARD OUT1 and feedback signal is high enough to assume that FSS and PZT controller are also not responsible for additional noise.
From the other hand MC2 SUSPOS measured by OSEMS shows good coherence with GUR1_X
That means that MC2 is indeed driven by seismic motion. In order to figure out if this is the case for MC1 and MC3, I rotated GUR2 by 45 degrees. When it will calm down, I'll measure coherence between OSEMs and seismic motion.